A solution contains 0.9% of Sodium Chloride. Calculate the number of particles that contribute to the osmotic pressure in 1 Liter of solution. (Mol wt: 58.5). Assume that the sodium chloride is completely soluble and each mole of NaCl dissociates in two ions (Na +& Cl) at this concentration.

  1. Calculate the quantity of sodium chloride present in 1 liter of solution.

  2. How many molecules are present in 1 liter?

  3. How many ions (particles) are formed when 1 mole of sodium chloride dissociates?

  4. How many ions (particles) are present in 0.9% sodium chloride?
Answer #1

The number of particles that contribute to the osmotic pressure in 1 Liter of solution = 1.20 x 1024

0.9 % means 0.9 grams of NaCl present in 100 mL of solution

a)

0.9 g / 100 mL ==> 9 grams in 1000 mL = 1 L

Answer: the quantity of sodium chloride present in 1 liter of solution = 9 g

b)

Avogadro's number NA = 6.02214129 x 1023 formula units/mol

moles NaCl in 9g = 9g / 58.5 g/mol = 0.15384615384 mol

molecules present in 1 L of solution = 0.15384615384 mol x 6.02214129 x 1023 molecules = 9.26 x 1022

Answer: molecules present in 1 L of solution = 9.26 x 1022

c)

Each molecule of NaCl dissociates into 2 ions NaCl ==> Na+ + Cl-

1 mole gives 2 ions, so ions formed when 1 mole of sodium chloride dissociates = 2 x 6.02214129 x 1023

Answer: ions (particles) are formed when 1 mole of sodium chloride dissociates = 1.20 x 1024

d)

moles in 0.9 g = 0.9g / 58.5 g/mol = 0.015384615384 mol

molecules present in 0.9% of solution = 0.015384615384 mol x 6.02214129 x 1023 molecules = 9.26 x 1021

ions formed = 2 x 0.015384615384 mol x 6.02214129 x 1023 molecules = 1.85 x 1022

Answer: ions (particles) are present in 0.9% sodium chloride = 1.85 x 1022

Hope this helped you!

Thank You So Much! Please Rate this answer as you wish.("Thumbs Up")